EDITORIAL

Terry Goodman

Another school year already! I can remember as a child hearing my grandparents say that the older you got the faster time passed. At the time I thought that sounded pretty silly. However, I have found that what they said was true!

The start of a new school year often causes us to reflect on teaching and learning mathematics. Perhaps we have not become so busy yet and we believe we can actually accomplish all we plan to do! Many of you are aware that the National Council of Teachers of Mathematics released, in April, its Principles and Standards for School Mathematics. This document “is intended to be a resource and guide for all who make decisions that affect the mathematics education of all students in prekindergarten through grade 12.” This document seeks to monitor and update the NCTM standards proposed in Curriculum and Evaluation Standards for School Mathematics (1989).

We all recognize that during the past decade there has been a great deal of discussion concerning mathematics education in this country. The impetus for much of that discussion came from the 1998 NCTM Standards. I think it is significant that NCTM has now produced a document that we can use to reflect on what has happened in the last decade as well as encourage us to think about the future of mathematics education.

I believe that the Principles and Standards can provide us with a valuable tool to use in the ongoing discussion concerning mathematics education in this country. It is significant that a number of professional organizations have already contributed to the development of the Principles and Standards. These groups included the American Mathematical Society, the Mathematical Association of America, and the American Mathematics Association of Two-Year Colleges. Certainly, there is a growing awareness among college/university faculty and many in the private sector that the mathematical understanding of K-12 students is of vital importance.

Congress and the various state legislatures often call for educational reform that in the end is superficial and greatly oversimplifies complex issues. The Principles and Standards represents an effort by NCTM and other professional organizations to be proactive in this discussion. I believe this document provides us with resources that we can use to give direction and focus to discussions about future directions for the mathematics education of students.

If you have not studied the Principles and Standards, I would encourage you to do so. You will find an electronic version of the document at http://www.nctm.org/standards. MJMS would welcome letters discussing issues related to the recommendations found in this document.

Once again, we would like to use the editorial pages in this issue to acknowledge those individuals who serve as referees for MJMS. A thorough and thoughtful peer– review process contributes greatly to the quality of a professional journal. Each of our referees plays a significant role in maintaining the quality of this journal and we deeply appreciate such consistent effort. If you are interested in serving as a referee for MJMS, we want to encourage you to complete the Referee Application Form found after the Announcements section in this issue.

List of Referees

Hadi Y. Alkahby, Dillard University Charles J. Allard, Polo R-VII High School Mohammad K. Azarian, University of Evansville Seung–Jin Bang, Seoul, Korea Gerald Bergum, South Dakota State University John E. Bernard, University of Texas–Pan American Jim Bruening, Southeast Missouri State University Joseph Chance, University of Texas–Pan American Hang Chen, Central Missouri State University Eungchun Cho, Kentucky State University Gregory Davis, University of Wisconsin–Green Bay Bryan Dawson, Union University Richard Delaware, University of Missouri–Kansas City Joseph B. Dence, University of Missouri–St. Louis Albert Dixon, College of the Ozarks Clayton Dodge, University of Maine John J. Edgell, Jr., Southwest Texas State University Vince Edmondson, Central Missouri State University Larry Eifler, University of Missouri–Kansas City Ekemezie J. Emeka, Columbia College Martin Erickson, Truman State University Russell Euler, Northwest Missouri State University Joe Flowers, Trinity University Jayanthi Ganapathy, University of Wisconsin–Oshkosh Ramesh Garimella, Northwest Missouri State University Raymond J. Grinnell, University of the West Indies Victor Gummersheimer, Southeast Missouri State University K.L.D. Gunawardena, University of Wisconsin–Oshkosh

List of Referees (cont.)

Dennis R. Harmon, Missouri Southern State College Stan Hartzler, Virginia Polytech Institute & State University Troy Hicks, University of Missouri–Rolla Joe Howard, New Mexico Highlands University Timothy P. Keller, George Mason University N. J. Kuenzi, University of Wisconsin–Oshkosh Charles E. Lamb, University of Texas–Austin Mulatu Lemma, Savannah State University Man Chun Leung, National University of Singapore Yi–Hsin Liu, University of Nebraska–Omaha William Livingston, Missouri Southern State College Roger Marty, Cleveland State University Wayne McDaniel, University of Missouri–St. Louis Gary McDonald, Northwest Missouri State University Rhonda L. McKee, Central Missouri State University Phoebe McLaughlin, Central Missouri State University Peter R. Mercer, St. Mary’s College of Maryland Kandasamy Muthvel, University of Wisconsin–Oshkosh Ollie Nanyes, Bradley University Krzysztof Ostaszewski, University of Louisville Miguel Paredes, University of Texas–Pan American Don Redmond, Southern Illinois University Jeffrey A. Rosoff, Gustavus Adolphus College Prasanna Sahoo, University of Louisville Jau–Shyong Shiue, University of Nevada–Las Vegas Donald P. Skow, University of Texas–Pan American Randall J. Swift, Western Kentucky University Jinhua Tao, Central Missouri State University Songlin Tian, Central Missouri State University Mohan Tikoo, Southeast Missouri State University Mathew Timm, Bradley University Alvin Tinsley, Central Missouri State University Peter L. Vachuska, University of Wisconsin–Washington Co. Bobby Neal Winters, Pittsburg State University Fred Worth, Henderson State University Joseph Wiener, University of Texas–Pan American